OpenStudy (anonymous):

How many milligrams of salt must be added to 18 milligrams of a 20% salt solution to obtain a 35% salt solution? Round to the nearest tenth.

OpenStudy (anonymous):

20 % = mass salt x 100/ 18 mass salt = 20 x 18/100=3.6 mg 35 = ( 3.6 + x) 100/ 18 + x 630 + 35 x = 360 + 100 x 270 = 65 x x =4.2 g